### v3

# Diagrammatic expansion for the mutual-information rate in the realm of limited statistics

Abstract
Neurons in sensory systems encode stimulus information into their stochastic spiking response. The mutual information has been extensively applied to these systems to quantify the neurons’ capacity of transmitting such information. Yet, while for discrete stimuli, like flashed images or single tones, its computation is straightforward, for dynamical stimuli it is necessary to compute a (mutual) information rate (MIR), therefore integrating over the multiple temporal correlations which characterize sensory systems. Previous methods are based on extensive sampling of the neuronal response, require large amounts of data and are therefore prone to biases and inaccuracy. Here, we develop Moba-MIRA (moment-based mutual-information-rate approximation), a computational method to estimate the mutual information rate. To derive Moba-MIRA, we use Feynman diagrams to expand the mutual information to arbitrary order in the correlations around the corresponding value for the empirical spike count distributions of single bins. As a result, only the empirical estimation of the pairwise correlations between time bins and the single-bin entropies are required, without the need for the whole joint probability distributions. We tested Moba-MIRA on synthetic data generated with generalized linear models, and showed that it requires only a few tens of stimulus repetitions to provide an accurate estimate of the information rate. Finally, we applied it to ex-vivo electrophysiological recordings of rats retina, obtaining rates ranging between 5 to 20 bits per second, consistent with earlier estimates.
I Introduction
Neurons in sensory systems respond to stimulus presentation by changes in their electrical potential, often with the emission of spikes, which are then sent to downstream areas for further processing [1]. Accordingly, stimulus information is encoded in the timing and frequency of those spikes [2, 3, 4, 5, 6, 7]. This information transmission is canonically estimated by the mutual information between the stimulus and the neuronal spiking response [8, 5]. In case of discrete and static stimuli this is done by counting the number of spikes emitted in a temporal window of a few hundreds of milliseconds following the stimulation. From the distribution of those spike counts it is then possible to estimate their empirical entropies and therefore the mutual information. In many applications, however, input and output are dynamical, information is transmitted  [^1] From a mathematical point of view, this word is a bit imprecise because it implies a directionality not reflected in the definition of the mutual information, which is symmetric in in- and output. Therefore, “shared” would be a more neutral way to express this relation. However, in the practical applications we have in mind, involving a stimulus as input and a neural activity as output, this directionality is given by the setup, so we will also use the term “transmitted”. over a certain period of time [6, 7] and the static method for estimating mutual information is inappropriate. In this case, it makes sense to consider the mutual information per time - the mutual-information rate (MIR [10]) - or the mutual information per emitted spike [11, 7, 12, 13, 14, 15]. Being able to reliably quantify this information is a necessary step to develop a quantitative understanding of sensory processing.
In order to achieve the goal of estimating the MIR, the data-intense histogram method to determine the entropies of long spiking patterns has been applied in several works [7, 12, 13, 14, 15]. Because this approach, also known as direct method, suffers in the data-limited case, a number of techniques have been developed to regularize the estimation and correct biases [16, 7]. These improvements have extended the range of applicability of the histogram method, which is however still limited to cases with relatively large datasets. Recently, Mahuas et al. [17] have proposed a complementary approach that avoids histograms developing MIR as a series in the empirical correlations. As this approach requires binarized neurons, it can deal only with small time bins, which limits its applicability.
In this work we introduce the moment-based mutual-information-rate approximation (Moba-MIRA), a method for estimating MIR from noisy data generalizing the work of [17]. Our method grounds on a diagrammatic expansion of the entropies in term of correlations. Recent results in field theory [18, 19, 20] allow us to expand around any non-interacting theory, like Ising spins as in [17], but also the empirical-spike-count (integer) distribution, and solving the issue occurring of small time bins.
We apply our method on data from the retina, a part of the nervous system that is particularly adapted to being studied by means of information theory because its input is very well controlled and it is relatively well accessible for recordings of neural activity. In the retina, incoming light is first absorbed by photoreceptors, transformed into an electrical signal, processed by a sequence of neurons and eventually encoded by the spiking activity of retinal ganglion cells (RGC) [21]. It has been attempted in several studies to estimate the resulting $\mathrm{MIR}$ by direct methods and Gaussian approximations [11, 22, 14, 23, 15, 24], yet, as indicated before, the approaches in these works were data intensive and have therefore a limited range of applicability. While we use our technique for the retina as a handy model system, it is much more widely applicable. One can use it not only for other neural systems, but virtually any system, not necessarily biological, whose behavior can be described as the response in form of a stochastic integer number to some input.
After briefly discussing the definition of the mutual-information rate in II, we derive Moba-MIRA in III. We first test its accuracy, in III.2, by applying it to synthetic data for which the ground truth value of the $\mathrm{MIR}$ can be estimated. In IV, we perform an in-depth analysis of retinal recordings, across stimuli and cell types and conclude in V, giving an outlook on possible further developments.
II Background
How does one compute the rate of transmitted information? A naïve approach to the problem, that follows from the static case, is to define MIR as
[…]
that is, the mutual information between stimulus and spike count, binned with a bin size $\Delta t$, and divided by $\Delta t$, where
${\cal I}\left(\Delta t\right)=S^{\mathrm{out}}\left(\Delta t\right)-\left\langle S_{\xi}^{\mathrm{in}}\left(\Delta t\right)\right\rangle,$ (2)
with $S^{\mathrm{out}}\left(\Delta t\right)$ the output (or marginal) entropy and the input (or conditional) entropy $S_{\xi}^{\mathrm{in}}\left(\Delta t\right)$ conditional on the stimulus. The choice of $\Delta t$ should depend on the system dynamics, and in particular on the relevant time scale of the stimulus. Yet, the dynamics of biological systems as the retina extends over multiple time scales, and cannot be captured by a single time bin. In order to understand the consequences of this choice, we consider the toy example of a neuron firing according to an inhomogeneous Poisson process. Mimicking the effect of a dynamical stimulus, the neuron’s rate randomly switches between a low and high state with an average frequency of $27\,\mathrm{Hz}$, cf. 1a. This yields an exponential decay of the autocorrelation of the neuron, both for the mean activity over repetitions (peristimulus time histogram, PSTH) and for its spike count (respectively, blue dashed and black line in 1b). Using (LABEL:Def_MIR_naiv) to compute the MIR for this process leads to an estimate monotonously decaying with the time-bin size (1c), showing that the choice of the bin duration strongly affects the MIR estimation. For large $\Delta t$ the stimulus dynamics is averaged out, and the MIR vanish. Upon decreasing $\Delta t$ this effect reduces, with the estimate for the mutual-information rate monotonously attaining a limiting value for $\Delta t\rightarrow 0$. However, this is not a generic behavior, as we figure out by investigating the same system, but after adding a refractory period to the model neuron, i.e. it stays silent for $10\,\mathrm{ms}$ after every spike, cf. 1d. The other parameters are unchanged. This leads to lower firing rates, more regular spike trains and spike-count autocorrelations which are negative and large for small times (1e). Refractory periods decrease the neurons’ variability, therefore increasing their capacity of transmitting information at fixed firing rate [25]. In order for the effect of the refractory period to become noticeable in the $\mathrm{MIR}$, however, $\Delta t$ has to be large enough (for a more detailed explanation see A.2 in the appendix). Consistently, the naïve MIR is not monotonic anymore and shows a maximum at around $15\,\mathrm{ms}$ (1f). While in the limit of $\Delta t\rightarrow 0$ we avoid stimulus averaging, we will neglect the positive effect of refractoriness, leading to an underestimation of the MIR. So, already in the case with only two time scales, there is no good choice for the length of $\Delta t$ fully accounting for the system dynamics.
Figure 1: (a) A neuron emitting spikes according to a rate switching randomly (Poisson process) between two levels, whereas the spiking conditional on the rate is either Poissonian itself (also panels b and c) or (d) features a refractory period (also panels e and f), a dot indicates a spike. (b,e) Autocorrelations of the mean activity indicated by the dotted blue lines, autocorrelations of the spikes by the solid black lines. (c,f) Mutual-information rates computed lumping activity into one time bin, as in (LABEL:Def_MIR_naiv), as a function of the time-bin size $\Delta t$. Parameters: Correlation time of switching of the rate: $T_{\mathrm{switch}}=100\,\mathrm{ms}$, firing rate $f_{\mathrm{low}}=5\mathrm{Hz}$, $f_{\mathrm{high}}=50\mathrm{Hz}$, (absolute) refractory period $t_{\mathrm{ref.}}=10\,\mathrm{ms}$, simulation time $T_{\mathrm{total}}=3\cdot 10^{4}\,\mathrm{ms}$, $N_{\mathrm{rep.}}=10^{4}$.
In order to solve the problem of multiple time scales, it was proposed in previous works to compute the MIR as [10, 11, 12]:
$\mathrm{MIR}\coloneqq\lim_{\mathrm{dt}\to 0}\underset{\Delta t\rightarrow\infty}{\lim}\frac{{\cal I}\left(\mathrm{dt},\Delta t\right)}{\Delta t},$ (3)
where the mutual information is computed over a large temporal window $\Delta t$, but after binning the neuron’s activity in small consecutive bins $\mathrm{dt}$, 2. Subdividing $\Delta t$ in $k$ time bins, as sketched in 2a, we obtain the estimates of the $\mathrm{MIR}$ for the model introduced in 1 and plot it in 2b,c as functions of $\Delta t$. We observe that the estimates converge to constant, non-zero values for $\Delta t\rightarrow\infty$. The dependence of the $\mathrm{MIR}$ estimate on $\Delta t$ can be explained as follows: the estimates of $\mathrm{MIR}$ initially decrease because information is encoded (slightly) redundantly in the time bins, so that adding more of them increases the mutual information (slightly) sublinearly. For large $\Delta t$ the activity in the newly added time bins is sufficiently distant from the first so that adding more time bins does not change the estimate of $\mathrm{MIR}$: we reach convergence. For the process including a refractory period, we have the same behavior for large $\Delta t$. For small $\Delta t$, we observe an initial increase with $\Delta t$ of the estimate of the $\mathrm{MIR}$ (panel c), which is absent without refractoriness. Qualitatively this is the same behavior as of the naive estimate of the $\mathrm{MIR}$ shown in 1. It can be explained by the fact that the activity, which is positively correlated between adjacent time bins for small $\Delta t$, gets decorrelated by the refractoriness. More precisely, the refractory period leads to negative noise (auto-)correlations canceling the positive stimulus (auto-)correlations. We discuss this effect in more detail in A.2. In the Poisson process without refractory period (panel b), the spike counts can be very high already for $\mathrm{dt}=10\,\mathrm{ms}$ and therefore, computing the $\mathrm{MIR}$ for a large value for $k$ becomes infeasible. We therefore have chosen $\mathrm{dt}=20\,\mathrm{ms}$ for panel c.
Figure 2: (a) Sketch of a recording of spiking neurons responding to a repeated stimulus over a time span $\Delta t$ with the activity discretised into bins of length $\mathrm{dt}$. (b) estimates of the $\mathrm{MIR}$ as (LABEL:MIR_def_proper), computed by the (mixed) Moba-MIRA method, introduced in III, for varying recording time $\Delta t$, along with their naive estimates, (LABEL:Def_MIR_naiv), as shown in 1, panel c. Time-bin size $\mathrm{dt}=20\mathrm{ms}$. (c) As panel b, but with refractory period, as in 1f. Other parameters as in 1.
The required entropies for the $\mathrm{MIR}$s as shown in 2 are those of probability distributions of stochastic paths across multiple time bins. They are high-dimensional objects because $\Delta t$ has to be large enough to cover the correlation time of the stimulus (and the response) and $\mathrm{dt}$ has to be small enough to capture the dynamics of the neural activity. For a limited amount of data, as is typical for experiments, these entropies quickly become difficult or impossible to compute reliably. The direct method to determine them, namely counting the occurrence of patterns (histogram method [7]) can therefore only work in simple setups, in particular for single neurons [6, 12, 13, 14, 15]. Even then, the number of time bins is limited (up to eight in the cited examples) because the number of possible words grows exponentially in this quantity. This issue applies in particular to the input entropy (conditional on the stimulus, cf. (LABEL:Def_MutualInfo)), because it quantifies the variability given a stimulus. It can therefore only be measured by repeating the same stimulus several times and collecting statistics over these repetitions. Here, therefore, number of samples equals that of the repetitions of the same stimulus, so typically at most about 100 in real data. In contrast, the output entropy marginalizing over the stimulus is computed over all times and repetitions, therefore the underlying probability distribution is much better sampled.
Even though there are sophisticated methods to improve the histogram method [7, 26, 27, 28, 29, 30], it is based on the characterization of a very high-dimensional probability distribution corresponding to an exponential number of moments. This is different if the data is Gaussian, which means that it is completely characterized by its first two moments. In this case there are closed-form expressions for the corresponding entropies vastly simplifying its computation [8, 31, 32, 33]. However, this prerequisite is often not met. Another option to regularize computations for the entropies consists in assuming a model for the underlying stochastic process, which can be used to parameterize the corresponding probability distributions. This is the case, for example, for chemical reaction networks, for which a host of new methods to compute the $\mathrm{MIR}$ have been developed in recent years [34, 35, 36]. However, such a model is not available in all situations and even if it is, it has to be fitted to the data, which can be a non-trivial - or at least numerically expensive - step on its own.
III Results: Moba-MIRA, a robust method to compute the MIR
III.1 Maximum-entropy modeling
Estimating entropies for real data is difficult because it provides only a limited number of repetitions, which leads to large biases. Our suggestion to mitigate this problem is to compute the entropies of the activities in the single time bins and to approximate the impact of correlations between them only on a pairwise level. Performing an expansion in the corresponding pairwise correlations and partly resumming this series leads to the following approximation for the entropy:
$S\approx S_{0}+\frac{1}{2}\left(\ln\left(\det\left(c\right)\right)-\ln\left(\det\left(V\right)\right)\right),$ (4)
where $S_{0}$ is the entropy of all single bins summed up, neglecting correlations, $c$ is the covariance matrix between bins (the autocovariance of the neuron under scrutiny across time) and $V$ are the respective variances, written as the entries of a diagonal matrix. This is what we call the moment-based mutual-information-rate approximation (Moba-MIRA). More precisely, we will use two versions of it: in one, we will use (LABEL:Entropy_single_plus_loops) only for the input entropy, while estimating the output entropy by the histogram method. We will call this variant the mixed Moba-MIRA, whereas we christen full Moba-MIRA the variant for which we use (LABEL:Entropy_single_plus_loops) for both types of entropies.
The approximation of (LABEL:Entropy_single_plus_loops) makes sense intuitively: it is nearly the Gaussian approximation, but with the important difference that we exactly take into account the entropies of the activitities (not necessarily binary) in single time bins. In the following we derive (LABEL:Entropy_single_plus_loops) in more grounded way using results from statistical physics and a diagrammatic expansion and explain how to add additional higher-order corrections.
III.1.1 Derivation of the Moba-MIRA approximation
Given a vector $\boldsymbol{n}=\left(n_{1},\dots,n_{k}\right)$ of spike counts recorded for the duration of $k=\frac{\Delta t}{\mathrm{dt}}$ time bins, depending on some other variable (e.g. a stimulus), which is identically repeated $N_{\mathrm{rep}}$ times, we want to estimate the probability distribution $P\left(\boldsymbol{n}\right)$ and the corresponding entropy $S$. We formalize the separation between statistics of the single bins, which we treat exactly, and the correlations between then, which we treat on a pairwise level, by making the following ansatz:
$P\left(\boldsymbol{n}\right)\sim e^{\frac{1}{2}\sum_{t\neq t^{\prime}}n_{t}J_{tt^{\prime}}n_{t}}\prod_{t=1}^{k}e^{-H_{t}\left(n_{t}\right)},\,\boldsymbol{n}\in\mathbb{N}^{k},$ (5)
where $H_{t}$ is some function and $\left\{J_{tt^{\prime}}\right\}_{1\leq t<t^{\prime}\leq k}$ is a matrix, which are both determined in order to match the measured statistics. Concretely, we choose
$H_{t}\left(n\right)=\sum_{i=1}^{\infty}\lambda_{i}n^{i},$ (6)
which is what results from a maximum-entropy modeling approach [37] with all moments of the single-bin statistics fixed. It is the formal expression of the informal statement that we treat the single-time-bin statistics exactly. At first sight, it appears that an infinity of quantities has to be measured. In practice however, this is not the case because the number of spikes observed per time bin is of course finite for finite recordings. Therefore there are at most $kn_{\mathrm{max}}$ quantities to determine for this contribution to the entropy, with $n_{\mathrm{max}}$ being the maximum number of spikes observed in one time bin. Together with the $\frac{k\left(k-1\right)}{2}$ correlations, this yields a number growing only quadratically with $k$. This is much better than the histogram method with its $\left(n_{\mathrm{max}}+1\right)^{k}$ parameters, which leads to considerable biases when $N_{\mathrm{rep}}=\mathcal{O}\left(10^{2}\right)$ - the relevant regime for experiments. Our method, in contrast, performs well there, as we will demonstrate later.
Our approach is the natural generalization of the traditional maximum-entropy framework employing binary variables [38], for which, as well, the single-unit statistics is reproduced (because fixing the mean already fixes the whole one-parameter distribution). Choosing binary variables, however, allows to take into account maximally one spike in every bin, otherwise information is lost. Sticking to this convention would therefore limit us in the choice of the time-bin width $\mathrm{dt}$. This would be unfortunate for our purposes, because we want to study the behavior of our estimate for the $\mathrm{MIR}$ in dependence on arbitrary $\mathrm{dt}$.
One possibility to use (LABEL:Def_generalized_Ising) would now be to fit the parameters $J_{tt^{\prime}}$ and in $H_{t}$ and then sample from $P$ to compute the entropy. However, we can get around this step by leveraging recently developed techniques from statistical field theory [18, 19, 20], explained in more detail in app. A.1. They allow us to derive the approximation (LABEL:Entropy_single_plus_loops) as a resummation of a class of diagrams in the diagrammatic small-correlation expansion around the case of uncorrelated time bins
{fmffile}
Gamma_loop_expansion
$S\approx$
which, using Feynman-diagrammatic rules  [^2] In statistical physics one sometimes also refers to them as Mayer diagrams [53, 54]. The work of Mayer & Goeppert-Mayer is actually a few years older than that of Feynman. indeed yields (LABEL:Entropy_single_plus_loops). Note that the elements of the diagrams represent combinations of the measured cumulants - we do not have to know $J_{tt^{\prime}}$ or $H_{t}$ explicitly. We will sketch the basics of diagrammatics in appendix A.1 and refer to [18, 19, 20] for a more detailed description. Considering corrections to (LABEL:Approx_SingleBinEnt_Loops) by taking into account more diagrams can be beneficial in some cases, but we have found the resummed-loop approximation to be the most robust.
III.2 Testing the approximation on artificial data
In order to validate MoBa-MIRA on a biologically plausible example, we use a generalized linear model (GLM) fitted to generate spike trains resembling those of retinal ganglion cells, for which we adapt the setup employed in [17] (see there and appendix A.3 for details). As shown in A1, when properly fitted, the GLM generates interspike-interval distributions and PSTHs barely distinguishable from real data. At the same time, when repeating a given stimulus to estimate the input entropies, we are only limited by the capacities of our computer. This allows us to compute a numerically exact value for biologically plausible data as ground truth.
In this setup, we can study in detail how the estimate of $\mathrm{MIR}$ changes as a function of $\mathrm{dt}$ and $\Delta t$. As visible from 3a, the estimate has converged at about $\Delta t=100\,\mathrm{ms}$. Fixing this value and varying $\mathrm{dt}$, we observe in 3b that we reach convergence at about $\mathrm{dt}=10-15\,\mathrm{ms}$. We therefore fix $\Delta t=100\,\mathrm{ms}$ and $\mathrm{dt}=10\,\mathrm{ms}$ for 3d, in which we plot the dependence of different estimates on the number of repetitions. Whereas the direct method shows a clearly visible bias at $N_{\mathrm{rep}}=\mathcal{O}\left(10^{2}\right)$, Moba-MIRA is already converged at about $N_{\mathrm{rep}}=50$. This is not a trivial consequence of negligible correlations between time bins as the comparison with the result neglecting correlations - that is, only summing up the entropies of single bins - shows.
How does the dependence of $\mathrm{MIR}$ estimates on $\mathrm{dt}$ and $\Delta t$ in the regime of limited data looks like for other methods? In 4a, we demonstrate that assuming the data to be simply Gaussian, which amounts to setting $S=\ln\left(\det\left(c\right)\right)+\mathrm{const.}$ leads to a clear overestimation of the $\mathrm{MIR}$ (yellow bar). Also, just neglecting correlations between time bins is not feasible, as the green bar shows, indicating that actually about half of the information is captured by the interaction between bins. Characterizing this relation using the histogram method, assuming a quadratic dependence on $1/N_{R}$ and extrapolating to $1/N_{R}=0$, as suggested by Strong et al. [7], yields much better results (orange bar), but leads still to an overestimate. However, assuming a probability distribution of the form (LABEL:Def_generalized_Ising) for the probability distribution conditional on the stimulus and applying the approximation (LABEL:Entropy_single_plus_loops) (mixed Moba-MIRA) yields an excellent fit, provided that we remove the bias by subtracting the estimates obtained from shuffled data, see A.5 for details. Imposing the form of (LABEL:Def_generalized_Ising) for the output entropy as well (full Moba-MIRA) yields a slightly worse fit, however, still performs better than the method from [7].
Why does our approach, (LABEL:Entropy_single_plus_loops), work? In general pairwise probability distributions, (LABEL:Def_generalized_Ising), might not be appropriate because the data could be also shaped by higher-order interactions in addition. However, for the input entropy, the most important part of the covariance is due to the refractory period of the neurons, as visible in the auto-correlation of the spikes, cf. A1. In other words, given the stimulus, the activity in two time bins is correlated mostly because a spike in one time bin suppresses a spike in another one  [^3] Note also that this interaction is not directed in time - whenever there is a spike at some point in time, one knows that there could not have been a spike both before and after this time. This effective suppression of spikes in neighbored time bins is an intrinsically two-point like interaction - at least as long as the time bin is not considerably shorter than the refractory period. It therefore does not come as a surprise that this approach works well for the input entropy. For the output entropy, however, we cannot make a similar argument and indeed, the pairwise approach clearly works worse in this case, in particular for smaller time bins.
Figure 3: Test of our method on data from a generalized linear model (GLM). (a) Scheme of a GLM: firing rates are computed depending on a visual stimulus, according to which spikes are generated in a random way (cf. app. A.3 for details). (b) Estimate of the $\mathrm{MIR}$ in dependence of $\Delta t$ for $\mathrm{dt}=10\,\mathrm{ms}$ and $N_{\mathrm{rep}}=3\cdot 10^{4}$. (c) Dependence of the $\mathrm{MIR}$ on the time-bin size $\mathrm{dt}$ for $\Delta t=100\,\mathrm{ms}$ and $N_{\mathrm{rep}}=3\cdot 10^{4}$. (d) Dependence of the $\mathrm{MIR}$ estimates on the number of repetitions $N_{\mathrm{rep}}$, $\mathrm{dt}=10\,\mathrm{ms}$, $\Delta t=100\,\mathrm{ms}$. All results were obtained with mixed Moba-MIRA, they are similar for full Moba-MIRA, compare 4a. For parameters of the GLM consult app. A.3.
Figure 4: Benchmark of our method. (a) Estimates of different methods for $N_{\mathrm{rep}}=80$, ground truth computed with histogram method with $N_{\mathrm{rep}}=3\cdot 10^{4}$. (b) Run time for the input entropy for the different methods is shown, $N_{\mathrm{rep}}=1000$. (c) Dependence of the $\mathrm{MIR}$ estimate on $N_{\mathrm{rep}}$, as panel d of 3, but including the approach from [17] and the histogram method with the same de-biasing procedure applied as for Moba-MIRA (shuffling, see A.5). (d) Dependence of the $\mathrm{MIR}$ estimate on the time bin $\mathrm{dt}$.
IV Results: Application of Moba-MIRA on retinal stimulus response
Figure 5: The $\mathrm{MIR}$ computed with (full) Moba-MIRA from ex-vivo retina recordings. (a) Sketch of the experimental setup. (b) Statistical measures of an example neuron - autocorrelation of the PSTH (gray) and the spikes (black), distribution of the interspike-intervals in the left inset, the PSTH for an example period in the right inset. (c) Scatter plot of the $\mathrm{MIR}$ against the firing rate. We indicate by a cross the neuron analyzed in more detail in A2a. (d) Mutual information per spike, same data as for panel c.
We now apply our method to data recorded in ex-vivo experiments on rat retinas [41]. To generate this data, the dissected retinas were stimulated by different light patterns and the activity of their output layer, containing the ganglion cells, was recorded by a multi-electrode array (MEA), see 5a. After spike-sorting [42] ganglion cells were clustered into functional types, resulting in two large populations of OFF and ON alpha cells [41]. Firstly, we started working on the former cells and determined the value of $\mathrm{dt}$ and $\Delta t$ from the autocorrelation of spikes and PSTHs (peristimulus time histogram, 5b): the intrinsic dynamics of the neurons (refractory period) is in the range of 10-20 milliseconds, whereas the correlation time due to the stimulus is in the range of hundred milliseconds.
Accordingly, we choose $\mathrm{dt}=10\mathrm{ms}$ and $\Delta t=200\mathrm{ms}$ to estimate the values shown in the panels c and d. A closer inspection of the dependence of the estimates of the $\mathrm{MIR}$ on these parameters, as done for the artificial data in 3c, shows that this choice indeed yields a sufficient resolution and a sufficiently large time window, cf app. A.4. Comparing the individual $\mathrm{MIR}$s with the firing rates, we observe a positive correlation (panel c). This relation, however, is sublinear so that the mutual information per spike decreases with the firing rate (panel d), in agreement with [15, their fig. (3)].
We then asked if our findings were consistent across experiments, stimuli and cell types. Indeed, repeating the same analysis for another retinal preparation driven by the same stimulus, we observed smaller firing rates and smaller $\mathrm{MIR}$ values. This is therefore consistent with the correlation between these two quantities. Also, the same behavior was observed when comparing the retinal response to two different visual stimulations - two randomly moving bars and checkerboards (6b). We observed the same behavior across population of ON and OFF cells (6c), so that, again in agreement to what has been observed in [15], there is no qualitative difference between the two types.
Figure 6: Comparison of the relation between firing rates and $\mathrm{MIR}$s for different modalities. (a) OFF cells recorded during stimulation with a checkerboard pattern in two different retinal preparations. (b) Same population of OFF cells subjected to different stimuli. We indicate by a cross the neuron analyzed in more detail in A2b. (c) Populations of ON and OFF cells in the same retinal preparation.
The results we have presented so far were obtained with the “full” version of Moba-MIRA, that is, we have used the approximation (LABEL:Entropy_single_plus_loops) for the input and the output entropies. However, in the analysis of the artificial data, we have seen that mixed Moba-MIRA yields results visibly closer to the ground truth, cf. 4a. Also, we have found single ON-cells for which full Moba-MIRA seems to deliver unreliable results, see A.5. Luckily, the estimates for the $\mathrm{MIR}$ of our real data typically already converge at values for $\Delta t$ as low as $80\mathrm{ms}$, see A2. In this case $k=8$ so that it is feasible to compute the estimate for $\mathrm{MIR}$ also in the mixed-Moba-MIRA version. This allows us to compare our results shown in 5 and 6, obtained by full Moba-MIRA, by their counterparts obtained by mixed Moba-MIRA, which we show in III.2. We observe that there are no qualitative differences between these two approaches.
Figure 7: Left panels) Full Moba-MIRA and Right panels) mixed Moba-MIRA. (a,b) Data from 5, panels c and panel d (inset). (c,d) Data from 6, panel b (e,f) Data from 6, panel c.
V Discussion
In this work we address the problem of estimating mutual information rates for noisy neurons responding to a dynamical stimulus. As observed in the past [7, 14], binning the neuron activity into large temporal windows fails because the dynamics of these systems have multiple time-scales, and no matter the bin length, information at longer or shorter scales is neglected (1). To solve this issue, short time bins have been used in previous works [7, 12, 13, 14, 15], and then long sequences of them were considered, so as to integrate over all the relevant time scales. These approaches are data hungry, and therefore very sensitive to undersampling. In contrast, our moment-based mutual-information-rate approximation (Moba-MIRA) is a method based on maximum-entropy assumptions relying on the estimate of the single-bin entropies and low-order statistics, which require much less measurements. As opposed to related approaches [43, 17], it is not limited to a binary representation of neural activity, therefore allowing for larger time bins and consequently achieves a compromise between resolution and the feasibility of computations. Based on a diagrammatic expansion and a resummation, it allows to robustly estimate entropies and the mutual-information rate.
The purpose of Moba-MIRA is to estimate conditional and marginal entropy of the spiking activity. In our approach we assume a pairwise maximum-entropy distribution of integer neurons and then determine the corresponding entropy by a small-correlation expansion, similar to the expansion for the Ising model of [44]. To derive this approximation, our framework benefits from recent developments in the field-theory for non-Gaussian theories [18, 19] that allow for using Feynman diagrams to compute corrections, potentially at all orders. We applied Moba-MIRA to synthetic datasets for which the ground truth is known, and benchmark it against previous proposed approaches. Moba-MIRA outperforms them, especially in the data-limited regime (3d).
Lastly we applied Moba-MIRA on rat retinal recordings, proving its capabilities in practical applications. We estimated the mutual information rate for rat retinal ganglion cells in response to checkerboard and randomly-moving bars movies. For one experiment for the checkerboard stimulus, we obtained rates between about $5\,\mathrm{bits}/s$ and $20\,\mathrm{bits}/s$, corresponding to $0.5\,\mathrm{bits}/\mathrm{spike}$ and $2.2\,\mathrm{bits}/\mathrm{spike}$ (mean equal to $1.1\pm 0.4\mathrm{bits}/\mathrm{spike}$, $n=30$ cells, compare 6c). Our estimates are in the range of previous results from the literature ($2.0\pm 0.7\,\mathrm{bits/\mathrm{spike}}$ and $2.1\pm 0.6\,\mathrm{bits/\mathrm{spike}}$ for brisk and sluggish cells in the rabbit, respectively [14], similar range for other cell types [15]), perhaps slightly lower. This deviation is mostly because our firing rates are a bit higher on average, but could also, to a smaller part, be due to the circumstance that the method by [7] tends to overestimate the $\mathrm{MIR}$, compare also 3 and figure A3.
Estimating mutual information rates is a relevant challenge in computational neuroscience of sensory systems. Previous methods are based on data-intensive histogram methods [12, 13, 14, 15], which have then been refined with additional extrapolation techniques [16, 7]. In order to compute information rates, we followed a different approach, and developed an approximation scheme that requires only the empirical estimation of several correlation matrices and that of the single-bin entropies, without needing the full probability distribution.
With our approach we extend and generalize [17]. First we consider integer spike counts, instead of binary, allowing for longer time bins and therefore less statistics to fit, without the need for clipping. Additionally, our theoretical scheme allows for using Feynman rules to compute corrections at potentially all orders. Note also that due to employing these approximations, we can compute the entropies directly from easy-to-measure quantities like covariances, without the need to fit a statistical model or to even define one. This fitting would be a step with numerically non-negligible costs, requiring to, e.g., iterate over several rounds of Monte-Carlo simulations to fix the correct values of the couplings, which we avoid.
Employing Moba-MIRA, we assume that the spike counts over multiple consecutive time bins follow a pairwise maximum-entropy distribution [37, 38]. These distributions have been proven effective in modeling of the activity of neural populations both for marginal [38, 45] and conditional [46, 43] distributions. As explained before, it is theoretically sound that a pairwise model works well for the probability distribution conditional on the input because its correlation structure is mostly determined by the refractory period after each spike of a neuron. Indeed, if the time-bin size is of the order of the refractory period, a change in the statistics of one time bin influences the statistics in the neighbouring bins, which is an intrinsically pairwise (even local) interaction. It is also symmetric because the occurrence of a spike in a certain time bin makes it equally less probable that another spike will occur after that and that a spike has occurred before.
Also, the pairwise model (LABEL:Def_generalized_Ising) provides even more potential to reduce numerical costs: By using the alternative definition
$\mathrm{MIR}_{\mathrm{diff}}\coloneqq\underset{\Delta t\rightarrow\infty}{\lim}\lim_{\mathrm{dt}\to 0}\frac{{\cal I}\left(\mathrm{dt},\Delta t+\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\Delta t\right)}{\mathrm{dt}}$ (7)
for the mutual-information rate, one can reduce the numerical cost from being quadratic in $k$ to linear. Indeed, with the definition (LABEL:MIR_def_proper), one has to compute the correlations of $\frac{k\left(k-1\right)}{2}$ pairs of bins, assembled from the $k$ time bins in the time window $\Delta t$. With the definition (LABEL:MIR_diff_Def), however, one only has to compute the correlation of the activity in the most advanced time bin, corresponding to the interval $\left(\Delta t,\Delta t+\mathrm{dt}\right]$, with the $k$ bins inside $\left[0,\Delta t\right]$, therefore only $k$ correlations. As noted before [11], both definitions for the MIR are equivalent when the correlations between time bins decay to 0. In practice, the achievable $\Delta t$ might not always be large enough to make both expressions agree, but even then, (LABEL:MIR_diff_Def) is an appealing alternative because it is closer to a differential definition of a rate or flux.
Still, however useful the hypothesis that pairwise couplings are sufficient to describe the interactions across time bins might be for computations, it does not hold true in general for both distributions, as a mixture of pairwise MaxEnt distributions does not belong to the model family itself. To reduce the possible impact of this uncontrolled assumption, we have proposed a variant of Moba-MIRA (mixed Moba-MIRA), for which we approximate only the conditional entropies, while performing extensive histogram count for computing the marginal entropy, where all the available data points can be used for one entropy estimate. Even if this comes with additional computational costs, we observed a neat improvement on the overall performance.
While our theoretical framework allows for computing corrections of higher order in the pairwise correlations, we did not observe an improvement of the performance in this case. A possible explanation is that by assuming a pairwise distribution, we are neglecting higher-order correlations, and these might have a larger impact than higher order terms in the expansion in pairwise correlations. Quantifying the relative impact of all different terms is difficult and would require an extension of our framework. In principle, however, our expansion around non-Gaussian, integer neurons allows for including higher-order correlations, and we will generalise Moba-MIRA to include them in the future. As indicated, this will be particularly interesting for the output entropy, for which the pairwise approximation is fair, but not optimal and actually sometimes fails qualitatively. While we can deal with this problem by employing the histogram method for the output entropy (that is, use mixed Moba-MIRA), a faster method for these cases is desirable.
In this work we applied Moba-MIRA to estimate mutual-information rates of individual neurons. Our method can however be extended to account for populations by modeling the correlation between different neurons at different times, going beyond the study of zero-time-lag correlations and their impact on information [47, 48, 17, 49]. Even if undersampling might be an issue there, we expect Moba-MIRA to be very useful, as methods based on histogram approaches would require an even larger amount of data, often beyond existing experimental datasets. With Moba-MIRA, however, only the estimation of correlations matrices is required, which reduces the necessary dataset size. We thus expect that reliable estimates can be given at least for pairs of neurons, analyzed with a temporal resolution comparable to that employed in this study. Currently, extending Moba-MIRA in that direction can be hindered by the lack of ground truth estimation for large populations, and because of this we leave it for future developments.
Acknowledgements.
We thank Kyle Bojanek and Olivier Marre and for insightful discussions and Filippo Castellani for providing the sketch of the experimental setup in 5. We acknowledge funding by ANR-21-CE37-0024 NatNetNoise, by IHU FOReSIGHT (ANR-18-IAHU-01) and by Sorbonne Center for Artificial Intelligence-Sorbonne University IDEX SUPER 11-IDEX0004. This work has been done within the framework of the PostGenAI@Paris project and it has benefitted from financial support by the Agence Nationale de la Recherche (ANR) with the reference ANR-23-IACL-0007. Our lab is part of the DIM C-BRAINS, funded by the Conseil Régional d’Ile-de-France.
Appendix A Technical details and additional analysis
A.1 Computing entropies by a diagrammatic small-correlation expansion around a theory with given statistics
For the statistics of a spike train discretized into $k$ different bins, in each of which there can be up to $n_{\mathrm{max}}$ spikes, fully characterizing the statistics means assigning a probability to each of the $\left(n_{\mathrm{max}}+1\right)^{k}$ states (sometimes called words [7]). For big $n_{\mathrm{max}}$ and, in particular, big $k$, the number of states becomes very large, which prohibits a reliable estimation given limited data.
Our way out we suggest in this manuscript is to use less demanding statistical measures, like the covariance between the activities of different bins to estimate the entropy. We therefore make the ansatz
$P\left(\boldsymbol{n}\right)=\frac{1}{{\cal Z}}e^{\frac{1}{2}\sum_{i\neq j}n_{i}J_{ij}n_{j}}\prod_{i=1}^{N}e^{-H_{i}\left(n_{i}\right)},$ (8)
where $H_{i}$ is some function to be inferred and
${\cal Z}=\sum_{\boldsymbol{n}}e^{\frac{1}{2}\sum_{i\neq j}n_{i}J_{ij}n_{j}}\prod_{i=1}^{N}e^{-H_{i}\left(n_{i}\right)}$
is the partition function, for the probability distribution of $\boldsymbol{n}$. The log-likelihood of this distribution is given by
${\cal L}=\frac{1}{2}\sum_{i\neq j}J_{ij}\left\langle n_{i}n_{j}\right\rangle_{P}-\sum_{i=1}^{N}\left\langle H_{i}\left(n_{i}\right)\right\rangle_{P}-\ln\left({\cal Z}\right),$
where by $\left\langle\dots\right\rangle_{P}$ we denote the average over the empirical distribution. The function $H_{i}$ is arbitrary in our technical framework - one might choose, e.g., $H_{i}\left(n\right)=-\ln\left(n!\right)-e^{h_{i}}$ in case one would like to expand around a Poissonian theory. We, instead, implement a maximum-entropy approach choosing it as power series according to (LABEL:Hamiltonian_infinitely_many_coeff). With this choice, we force our statistical model to reproduce all empirically measured cumulants for single bins. This is ensured by choosing the interaction matrix $J$ and the (infinitely many) parameters of $\boldsymbol{H}$, $\lambda_{1},\lambda_{2},\dots$ accordingly. Because $\ln\left({\cal Z}\right)$ is the cumulant-generating function, we can express this condition as
${\cal L}_{\mathrm{max}}=\sup_{J,\lambda_{1},\lambda_{2},\dots}\left(\frac{1}{2}\sum_{i\neq j}J_{ij}\left\langle n_{i}n_{j}\right\rangle_{P}-\sum_{i=1}^{N}\sum_{\alpha=1}^{\infty}\lambda_{\alpha}\left\langle n_{i}^{\alpha}\right\rangle_{P}-\ln\left({\cal Z}\right)\right),$
where $\left\langle\dots\right\rangle_{P}$ is the empirical average. Once the parameters are fixed such that the measured statistics are reproduced by the model, we can also write
${\cal L}_{\mathrm{max}}=\sup_{J,\lambda_{1},\lambda_{2},\dots}\left(\frac{1}{2}\sum_{i\neq j}J_{ij}\left\langle n_{i}n_{j}\right\rangle-\sum_{i=1}^{N}\sum_{\alpha=1}^{\infty}\lambda_{\alpha}\left\langle n_{i}^{\alpha}\right\rangle-\ln\left({\cal Z}\right)\right)=-S,$
where we denote by $\left\langle\dots\right\rangle$ the average with respect to the model. In words: the negative maximum log-likelihood equals the entropy of the statistical model and also the free energy at fixed covariances and fixed single-neuron statistics.
To compute it in practice, we perform an expansion in small covariances. We will use Feynman diagrams for it, which simplify this endeavor because the corresponding rules incorporate the structure of the terms in the series in a compact and elegant way, which comes about by the fact that the free energy is a Legendre transform [19, 20]. Also, they allow to identify contributions to the series according to certain topologies of the diagrams, which can partly be resummed. For our work, we employ the resummation of all loops, as in
{fmffile}
Loop_expansion
[…]
where we call the vector of variances $\boldsymbol{V}$, understand $\frac{c}{\boldsymbol{V}^{\mathrm{T}}\boldsymbol{V}}$ as element-wise product and denote the diagonal matrix filled with the entries of $\boldsymbol{V}$ as $V$. The infinite sum is represented by Feynman diagrams, composed of the elements
{fmffile}
Collection_diagram_elements
Every diagram is evaluated by multiplying the mathematical objects its components represent, summing over all indices and including a factor, which depends on the number of vertices (edges) and the symmetry of the diagram. For a description of the precise Feynman rules, we refer to [18, 19, 20]. Note that it is possible to write the resummation in this compact form, using matrix operations, because we perform the Legendre transform not only with respect to the covariances, the off-diagonal part of the covariance matrix, as in [19], but also with respect to the variances, its diagonal part (and all other single-bin cumlants). As explained in detail in [20, sec. 2.2.3], this lifts restrictions in the sums of the perturbation expansion, which would prevent writing them as simple matrix multiplications.
The rule to determine the estimation of the entropy for the whole system is therefore quite straight-forward: Compute the entropies of the single bins, sum them up, then add correcting terms, coming about by the correlations between the bins and expressed by diagrams. By construction, the cumulants in the small-correlation expansion translated from diagrams are then the empirical ones - because the unperturbed theory is the maximum-entropy single-bin model reproducing the single-bin statistics.
A.2 The effect of a refractory period on estimates of the mutual-information rate
We will demonstrate in this section that for spiking activity with a (hard) refractory period $t_{\mathrm{ref}}$, the estimate for the $\mathrm{MIR}$ initially grows with $\Delta t$, as visible in 2c.
Assume that $\Delta t<t_{\mathrm{ref}}$, so that we only have to consider a binary representation of the activity and consider $k=2$, so $\mathrm{dt}=\Delta t/2<t_{\mathrm{ref}}/2$. In this two bins, the firing rate can attain two different values, which we call $\lambda_{1}$ and $\lambda_{2}$. We assume that the averages of both rates are identical. We can then explicitly compute the $\mathrm{MIR}$ based on the activities $n_{1}$ and $n_{2}$ of the corresponding time bins. First, we obtain for the input entropy
$S_{\mathrm{in}}\left(\mathrm{dt},\mathrm{dt}\right)=$ $-\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\ln\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)$
$-\mathrm{dt}\left(\lambda_{1}\ln\left(\lambda_{1}\mathrm{dt}\right)+\lambda_{2}\ln\left(\lambda_{2}\mathrm{dt}\right)\right)$
and for the output entropy
$S_{\mathrm{out}}\left(\mathrm{dt},\mathrm{dt}\right)=$ $-\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\ln\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$
$-\mathrm{dt}\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)+\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\right),$
which then yields for the mutual information
${\cal I}\left(\mathrm{dt},\mathrm{dt}\right)=$ $-\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\ln\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$
$+\left\langle\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\ln\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}$
$-\mathrm{dt}\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)+\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\right)$
$+\mathrm{dt}\left(\left\langle\lambda_{1}\ln\left(\lambda_{1}\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}+\left\langle\lambda_{2}\ln\left(\lambda_{2}\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}\right).$ (12)
We expand the first two lines in $\mathrm{dt}$ to obtain
$-\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\ln\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$
$+\left\langle\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\ln\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}$
$=$ $-\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\left(-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}-\frac{1}{2}\left(\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)^{2}\right)$
$+\left\langle\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\left(-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}-\frac{1}{2}\left(\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)^{2}\right)\right\rangle_{\boldsymbol{\lambda}}+\mathcal{O}\left(\mathrm{dt}^{3}\right)$
$=$ $-\frac{1}{2}\left(\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)^{2}+\frac{1}{2}\left\langle\left(\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)^{2}\right\rangle_{\boldsymbol{\lambda}}+\mathcal{O}\left(\mathrm{dt}^{3}\right)$
[…]
$2\mathrm{dt}\left(\left\langle\lambda\ln\left(\lambda\right)\right\rangle_{\lambda}+\ln\left(\mathrm{dt}\right)\left\langle\lambda\right\rangle_{\lambda}-\left\langle\lambda\right\rangle_{\lambda}\ln\left(\left\langle\lambda\right\rangle_{\lambda}\right)-\ln\left(\mathrm{dt}\right)\left\langle\lambda\right\rangle_{\lambda}\right)$
$=$ $2\mathrm{dt}\left(\left\langle\lambda\ln\left(\lambda\right)\right\rangle_{\lambda}-\left\langle\lambda\right\rangle_{\lambda}\ln\left(\left\langle\lambda\right\rangle_{\lambda}\right)\right).$ (13)
So, in total, we obtain obtain for the mutual information of the pair of time bins
[…]
Similarly, we obtain for the mutual information of the single bins
[…]
Therefore, the estimate of the $\mathrm{MIR}$ are, respectively
[…]
the difference being
[…]
which corresponds to the difference between the $\mathrm{MIR}$ estimates for $k=2$ and $k=1$ (or $\Delta t=2\mathrm{dt}$ and $\Delta t=\mathrm{dt}$ in 2). If the rates in two consecutive time bins are positively correlated, computing the $\mathrm{MIR}$ for both of the bins at once increases the estimate, otherwise it decreases it. This observation reminds of what is known as sign rule in the study of noise correlations in populations of neurons [50]: if noise and stimulus correlations have opposite sign, the mutual information increases compared to the case without noise correlations. Indeed, refractoriness leads to negative noise correlations and positive rate correlations leads to an increase of the $\mathrm{MIR}$, whereas negative rate correlations lead to a decrease.
A.2.1 Increasing the time bin
We can also lump together the activity of the two time bins into a larger one, of size $2\mathrm{dt}$, instead of two consecutive time bins of size $\mathrm{dt}$, as in 1. This changes the above computation a bit. We then have
${\cal I}\left(2\mathrm{dt}\right)=$ $-\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\ln\left(1-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$
$+\left\langle\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\ln\left(1-\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}$
$-\mathrm{dt}\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$
$+\mathrm{dt}\left\langle\left(\lambda_{1}+\lambda_{2}\right)\ln\left(\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}.$ (15)
The first two lines agree with the two-bin case, so that we have
${\cal I}\left(2\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\mathrm{dt}\right)=$ $-\mathrm{dt}\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)$ (16)
$+\mathrm{dt}\left\langle\left(\lambda_{1}+\lambda_{2}\right)\ln\left(\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}$ (17)
$-\left[\mathrm{dt}\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)+\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\right)\right$. (18)
$\left.+\mathrm{dt}\left(\left\langle\lambda_{1}\ln\left(\lambda_{1}\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}+\left\langle\lambda_{2}\ln\left(\lambda_{2}\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}\right)\right].$ (19)
We assume in the following that the firing rates fluctuate weakly in the sense that
$\lambda_{i}=\lambda_{0}+\delta\lambda_{i},\ \frac{\delta\lambda_{i}}{\lambda_{0}}\ll 1,\ i\in\left\{1,2\right\}.$
We note that this condition is not met for the example we employ in the main text, for 2, for which the rate switches between two rates, of which one is one order of magnitude smaller than the other one. If we use the following considerations to interpret our observations, we therefore have to take them with a grain of salt, to say the least. We find them instructive nonetheless.
We can expand (LABEL:Diff_MI_longbin) and (LABEL:Diff_MI_longbin_II) in small $\delta\lambda$
$\mathrm{dt}\left[\left\langle\left(\lambda_{1}+\lambda_{2}\right)\ln\left(\left(\lambda_{1}+\lambda_{2}\right)\mathrm{dt}\right)\right\rangle_{\boldsymbol{\lambda}}-\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\ln\left(\left\langle\lambda_{1}+\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}\mathrm{dt}\right)\right]$
$=$ $\mathrm{dt}\left[\left\langle\left(2\lambda_{0}+\delta\lambda_{1}+\delta\lambda_{2}\right)\ln\left(2\lambda_{0}+\delta\lambda_{1}+\delta\lambda_{2}\right)\right\rangle_{\boldsymbol{\lambda}}-2\lambda_{0}\ln\left(2\lambda_{0}\right)\right]$
$=$ $\mathrm{dt}\left[\left\langle\left(2\lambda_{0}+\delta\lambda_{1}+\delta\lambda_{2}\right)\left[\ln\left(2\lambda_{0}\mathrm{dt}\right)+\ln\left(1+\frac{\delta\lambda_{1}+\delta\lambda_{2}}{2\lambda_{0}}\right)\right]\right\rangle_{\boldsymbol{\lambda}}-2\lambda_{0}\ln\left(2\lambda_{0}\mathrm{dt}\right)\right]$
$=$ $\mathrm{dt}\left\langle\left(\delta\lambda_{1}+\delta\lambda_{2}\right)\left[\left(\frac{\delta\lambda_{1}+\delta\lambda_{2}}{2\lambda_{0}}\right)-2\lambda_{0}\frac{1}{2}\left(\frac{\delta\lambda_{1}+\delta\lambda_{2}}{2\lambda_{0}}\right)^{2}\right]\right\rangle_{\boldsymbol{\lambda}}+\mathcal{O}\left(\left(\frac{\delta\lambda_{i}}{\lambda_{0}}\right)^{3}\right)$
$=$ $\mathrm{dt}\frac{1}{4\lambda_{0}}\left\langle\left(\delta\lambda_{1}+\delta\lambda_{2}\right)^{2}\right\rangle_{\boldsymbol{\lambda}}+\mathcal{O}\left(\left(\frac{\delta\lambda_{i}}{\lambda_{0}}\right)^{3}\right).$ (20)
The last two lines of the expression for ${\cal I}\left(2\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\mathrm{dt}\right)$, (LABEL:Diff_MI_doublebin) and (LABEL:Diff_MI_doublebin_II), yield
$-2\mathrm{dt}\left(\left\langle\lambda\ln\left(\lambda\,\mathrm{dt}\right)\right\rangle_{\lambda}-\left\langle\lambda\right\rangle_{\lambda}\ln\left(\left\langle\lambda\right\rangle_{\lambda}\mathrm{dt}\right)\right)$
$=$ $-\mathrm{dt}\frac{\left\langle\left(\delta\lambda\right)^{2}\right\rangle_{\lambda}}{\lambda_{0}}+\mathcal{O}\left(\left(\frac{\delta\lambda}{\lambda_{0}}\right)^{3}\right).$ (21)
We observe that (LABEL:Rewrite_part_of_MI2dt) and (LABEL:Rewrite_part_of_dtdt) agree for the case of perfect correlation, otherwise the estimate from the large time bin is lower than that for two small time bins. For the difference between the mutual informations in the large time bin and in the two small time bins, this means
${\cal I}\left(2\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\mathrm{dt}\right)$
$=$ $\mathrm{dt}\frac{1}{4\lambda_{0}}\left\langle\left(\delta\lambda_{1}+\delta\lambda_{2}\right)^{2}\right\rangle_{\boldsymbol{\lambda}}-\frac{\mathrm{dt}}{\lambda_{0}}\left(\left\langle\delta\lambda^{2}\right\rangle\right)$
$=$ $\frac{\mathrm{dt}}{2\lambda_{0}}\left(\left\langle\delta\lambda_{1}\delta\lambda_{2}\right\rangle-\left\langle\delta\lambda^{2}\right\rangle\right)$
In our toy model with the rate of the imhomogeneous Poisson process switching itself in a Poisson fashion, we can make this more precise. We determine the correlation of the rate fluctuations by taking into account that there are only two possibilities: either the rate does not switch from one time bin to the next - which is true with probability $\exp\left(-\mathrm{dt}/T_{\mathrm{switch}}\right)$; then the two rate fluctuations $\delta\lambda_{1}$ and $\delta\lambda_{2}$ are identical, or there is a switch; then the fluctuations have opposite sign, which happens with probability $1-\exp\left(-\mathrm{dt}/T_{\mathrm{switch}}\right)$. Therefore, we have
$\left\langle\delta\lambda_{1}\delta\lambda_{2}\right\rangle_{\boldsymbol{\lambda}}$ $=\left\langle\left(\delta\lambda\right)^{2}\right\rangle_{\lambda}\left(2e^{-\frac{\mathrm{dt}}{T_{\mathrm{switch}}}}-1\right)$
$=\left\langle\left(\delta\lambda\right)^{2}\right\rangle_{\lambda}\left(1-2\frac{\mathrm{dt}}{T_{\mathrm{switch}}}\right)+\mathcal{O}\left(\mathrm{dt}^{2}\right)$
and thus
${\cal I}\left(2\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\mathrm{dt}\right)=-\frac{\mathrm{dt}^{2}}{\lambda_{0}T_{\mathrm{switch}}}\left\langle\delta\lambda^{2}\right\rangle+\mathcal{O}\left(\mathrm{dt}^{3}\right)$
and therefore
$\mathrm{MIR}\left(2\mathrm{dt}\right)-\mathrm{MIR}\left(\mathrm{dt},\mathrm{dt}\right)$ $=\frac{{\cal I}\left(2\mathrm{dt}\right)-{\cal I}\left(\mathrm{dt},\mathrm{dt}\right)}{2\mathrm{dt}}$
$=-\frac{\mathrm{dt}^{2}}{2\lambda_{0}T_{\mathrm{switch}}}\left\langle\delta\lambda^{2}\right\rangle+\mathcal{O}\left(\mathrm{dt}^{2}\right).$
[…]
$\mathrm{MIR}\left(2\mathrm{dt}\right)-\mathrm{MIR}\left(\mathrm{dt}\right)=\frac{1}{2}\mathrm{dt}\left\langle\delta\lambda^{2}\right\rangle\left(1-\frac{1}{\lambda_{0}T_{\mathrm{switch}}}\right)+\mathcal{O}\left(\mathrm{dt}^{2}\right).$
So if there is on average at least one spike between two switches of the firing rate, the estimate for the $\mathrm{MIR}$ will increase as well by increasing the time-bin size.
A.3 Adapting the generalized linear model
Building on [17], we adapt the parameters of the generalized linear model (GLM) in order to generate data resembling the one from experiments. To be precise, we compare the autocorrelations and the interspike-interval distributions (ISIs) of the artificial neuron to the corresponding measures from that cell in the the off population presented in the main text in 5c,d, stimulated by a black-and-white checkerboards. This neuron is quite typical for the population and seems to be a good example in terms of data quality, see A1. We refer to the supplemental material of [17] for a detailed description of the GLM and paraphrase here only the parts needed to understand our parameter changes.
The spiking rate in the GLM is here given by
$\lambda_{i}\left(t\right)=\frac{1}{1+e^{-h_{i}\left(t\right)}},\ h_{i}\left(t\right)=h_{i}^{\mathrm{bias}}+h_{i}^{\mathrm{stim}}\left(t\right)+h_{i}^{\mathrm{int}}\left(t\right),$
where $h_{i}^{\mathrm{stim}}\left(t\right)$ contributes the effect of the stimulus, $h_{i}^{\mathrm{int}}\left(t\right)$ the self-coupling of the neuron (mimicking refractoriness) and through $h_{i}^{\mathrm{bias}}$ one controls the basic level activity. $h_{i}^{\mathrm{stim}}\left(t\right)$ is given by the (temporal) convolution of the input with the difference of two raised cosine functions, as in
$\mathrm{rc}\left(\tau,s,c\right)=\begin{cases}\cos^{2}\left(\frac{\pi}{2}\left(\ln\left(\tau+s\right)-c\right)\right),&-1\leq\ln\left(\tau+s\right)-c\leq 1\\ 0,&\mathrm{otherwise}.\end{cases}$
To make the response of the neuron faster, we change the parameters of [17] according to $c_{1}=4.8\rightarrow 4.1$, $c_{2}=5.3\rightarrow 4.6$ and $s=50\rightarrow 25$, which compresses the kernel by approximately a factor $2$ (note the exponential relation between $\tau$ and $c$). Furthermore, we changed the overall prefactor entering in $h_{i}^{\mathrm{stim}}\left(t\right)$ from $0.5$ to $2$. Additionally, we change the bias from 4-4 to 3-3, which increases the mean firing rate. Finally, we replace the absolute refractory period of $10\,\mathrm{ms}$ by an absolute refractory period of $5\,\mathrm{ms}$, followed by a relative refractory period with exponential recovery with a decay time of $10\,\mathrm{ms}$. This yields the statistics shown in figure 7 (left column).
Figure A1: Comparison of the statistics of artificial and real data. We analyze the spiking activity of the model described in this to the recordings presented in IV, artificial data in the left, real data in the right column. The autocorrelation of the PSTH and the spikes are shown in panels a and b, respective insets show example section of the PSTH. Inter-spike intervals are shown in panel c and d.
A.4 Computing MIRs for real data in praxis
Figure A2: (a) The dependence of the $\mathrm{MIR}$ estimate on the time-bin size for the neuron of 5. Final estimate of the $\mathrm{MIR}$ (with $\mathrm{dt}$ fixed) indicated by blue cross. (d) Like panel a, but the experiment from 6.
In this part of the appendix, we explain how to choose appropriate values for $\mathrm{dt}$ and $\Delta t$ to analyze real data. To achieve this, we compare the estimate of the $\mathrm{MIR}$ for different values of $\Delta t$ and varying time-bin size $\mathrm{dt}$ in A2. The differently nuanced curves indicate different values of $\Delta t$. Because in our examples they nearly lie on top of each other, we are confident that our values $\Delta t$ with which we perform the final estimate of the $\mathrm{MIR}$ is large enough.
From the $\mathrm{dt}$-$\mathrm{MIR}$ curve, we can read of what is a good value for the time bin. We know that it should be decreasing because a lower time resolution decreases the information. For the neuron whose data is shown in A2a, this is the case and we therefore take the lowest value for $\mathrm{dt}$ for which convergence is about to be reached. In this experiment, there are 79 repetitions, quite a lot for a neuroscience experiment, whereas for experiment whose data is shown A2b, there are only 54 repetition, which makes it more challenging to analyze. Indeed, we observe that the corresponding $\mathrm{dt}$-$\mathrm{MIR}$ curve is not monotonous. This behavior might derive from the bias due to the lower number or repetitions. Also, it could be that our approximation in the regime of small time bins is imprecise for this neuron and small time bins.
By using A2, we can detect this behavior with our analysis and deal with it. As the best proxy, we choose the maximum of the curve. By making this choice, we assume that the true value of $\mathrm{MIR}$ at $\mathrm{dt}\rightarrow 0$ does not deviate much from its equivalent estimate at moderate time-bin size. Note that this is an assumption that, by its very nature, we cannot check because it would require to either reduce $\Delta t$, which is prohibitive because of the correlation time of the stimulus, or to increase the number of repetitions, which we cannot achieve either, of course. For some neurons we checked, the failure of full Moba-MIRA is even clearer than in the example shown in A2, in one case even leading to negative estimates in the limit of small $\mathrm{dt}$, shown in figure A3b. Note however, that for all the retinal recordings we checked the variant of mixed Moba-MIRA yields reasonable results - without a comparable failure. While it limits the range of number of time bins (and therefore $\Delta t$) that one can use, because the histogram method is numerically expensive, this is therefore still a resort that one can take, given that for our data no huge $\Delta t$ are necessary. An important use of plots like in A2 is to check that the approximation leads to reasonable results.
A.5 Strategies of removing the bias in estimates
How do we deal with the problem of biases in the estimation of entropies for high-dimensional data? Indeed, because the entropy is a concave function of the probability, an imprecise estimate of the latter leads to underestimating the entropy. Because we have many samples for the output entropy, the effect of fluctuations is negligible there, whereas for the input entropy, we only have as many samples as repetitions of the same stimulus, so at most about $100$, and therefore a much bigger bias. In the following, we will therefore focus on how to remove the bias from the input entropy.
Although the bias is reduced by projecting the measured statistics on a statistical model, it is still large enough to spoil the results for the typical number of repetitions available in recordings. To reduce this remaining bias, which is due to suboptimal estimates for the covariances across time bins, we employ what is known as shuffling approach [51, 17]: we shuffle the spike times across repetitions, removing the noise correlations. Conditional on the stimulus, the activity across time bins should therefore be independent and the estimate for the entropy equal that one for the single time bins. However, this will not be the case, precisely because of the bias explained before. Assuming that this bias is the same for the shuffled and the unshuffled data, we take the difference of the resulting entropy from the one of the single-bin estimate to estimate it. Concretely, we take
$S_{\mathrm{across\ bin},\mathrm{de-biased}}^{\mathrm{in}}=S_{\mathrm{across\ bin}}^{\mathrm{in}}-S_{\mathrm{across\ bin,\mathrm{shuffled}}}^{\mathrm{in}}+S_{\mathrm{single}}^{\mathrm{in}}.$
In addition, we also reduce the bias in the contribution from the single bins in a similar way: by shuffling the spike times across time, we make all time bins statistically equivalent, removing the variability due to the stimulus. In the shuffled data, each time bin therefore has the same statistics as the the neural activity sampled across all time bins. We therefore have
$\lim_{N_{\mathrm{rep}}\rightarrow\infty}S_{\mathrm{single},\mathrm{shuffled\ in\ time}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)=\lim_{N_{\mathrm{rep}}\rightarrow\infty}S_{\mathrm{single}}^{\mathrm{out}}\left(N_{\mathrm{rep}}\right).$
For a finite number of repetitions, however, also $S_{\mathrm{single},\mathrm{shuffled\ in\ time}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)$ will be biased. As before, we assume that this bias is the same as for the unshuffled version and therefore take
$S_{\mathrm{single},\mathrm{de-biased}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)=S_{\mathrm{single}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)-S_{\mathrm{single},\mathrm{shuffled\ in\ time}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)+S_{\mathrm{single}}^{\mathrm{out}}\left(N_{\mathrm{rep}}\right).$
This means for the mutual information that we have
${\cal I}_{\mathrm{single,\mathrm{de-biased}}}\left(N_{\mathrm{rep}}\right)=S_{\mathrm{single},\mathrm{shuffled\ in\ time}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right)-S_{\mathrm{single}}^{\mathrm{in}}\left(N_{\mathrm{rep}}\right).$
From figure A3, it is apparent that these additional steps are important for the good performance of our method, whereas they do not improve, but rather deteriorate, the estimate according to [7], based on the histogram method. We reckon that this comes about by the introduction of spurious higher-order correlations by the shuffling, to which the histogram method is sensitive. They lead to an overestimate of the bias, rendering the shuffling trick useless for this method.
In addition to the shuffling trick, we also regularize our estimates of the covariances, by taking
$c_{t}^{\mathrm{noise},\,\mathrm{est}}=\left(1-\epsilon\right)c_{t}^{\mathrm{noise}}+\epsilon\bar{c}^{\mathrm{noise}},$
where $\bar{c}^{\mathrm{noise}}$ is the time average over all noise covariances. This is a common procedure to reduce noise in this estimate and therefore the bias in the entropy [52].
Figure A3: (a) Dependence of the estimate from mixed Moba-MIRA on $N_{\mathrm{rep}}$ for the artificial data of 3 and 4, $k=10$. (b) Same for the histogram method [7], together with quadratic extrapolation in $\frac{1}{N_{\mathrm{rep}}}$ (fitted for all $N_{\mathrm{rep}}\geq 20$). (c) Estimate of $\mathrm{MIR}$ of an ON cell for different methods from the population presented in 6, panel c. (d) Same for off cell from the population shown in 5, panel c and d.